Write the equation of the line L satisfying the given geometric conditions.
(36) L has y-intercept (0, -3) and is parallel to the line with equation
y = 2/3x + 1.
(40) L has y-intercept (0,2) and is perpendicular to the line with equation 2x-3y = 6.

Answer by venugopalramana(3286)

Write the equation of the line L satisfying the given geometric conditions.
(36) L has y-intercept (0, -3) and is parallel to the line with equation
y = 2/3x + 1.
EQN.OF PARALLEL LINE IS OF THE FORM
Y =2X/3 +K.............IT PASSES THROUGH (0,-3).HENCE
-3=K
SO EQN. IS
Y=2X/3 -3
(40) L has y-intercept (0,2) and is perpendicular to the line with equation 2x-3y = 6.
EQN.OF PERPENDICULAR LINE IS OF THE FORM
3X+2Y=K.............IT PASSES THROUGH (0,2).HENCE
2*2=K=4
SO EQN.IS
3X+2Y=4